OpenStudy (anonymous):

I sold an article for $60 and made a profit of 20% what was the cost price?

OpenStudy (anonymous):

the cost price is 50

OpenStudy (anonymous):

Use this formula: percentage profit= new amount - original amount / original amount multiplied by 100. Let the percentage profit be 20 and the new amount is 60 but the original amount make it x. so u have: 20=60-x/x *100 20x=100(60-x) 20x=6000-100x 120x=6000 x=6000/120 x=50 so the cost price is $50 and to be sure substitute x with 50

OpenStudy (anonymous):

Thanks a whole lot Faris

OpenStudy (anonymous):

u r welcome